OpenStudy (anonymous):

Solve: (4x - 5)(3x + 2) =0

OpenStudy (anonymous):

\[ \left\{\left\{x\to -\frac{2}{3}\right\},\left\{x\to \frac{5}{4}\right\}\right\} \]